# Quality Diversity through Human Feedback

### v2

@@ Abstract
Reinforcement Learning from Human Feedback (RLHF) has shown potential in qualitative tasks where clear objectives are lacking. However, its effectiveness is not fully realized when it is conceptualized merely as a tool to optimize average human preferences, especially in generative tasks that demand diverse model responses. Meanwhile, Quality Diversity (QD) algorithms excel at identifying diverse and high-quality solutions but often rely on manually crafted diversity metrics. This paper introduces Quality Diversity through Human Feedback (QDHF), a novel approach integrating human feedback into the QD framework. QDHF infers diversity metrics from human judgments of similarity among solutions, thereby enhancing the applicability and effectiveness of QD algorithms. Our empirical studies show that QDHF significantly outperforms state-of-the-art methods in automatic diversity discovery and matches the efficacy of using manually crafted metrics for QD on standard benchmarks in robotics and reinforcement learning. Notably, in a latent space illumination task, QDHF substantially enhances the diversity in images generated by a diffusion model and was more favorably received in user studies. We conclude by analyzing QDHF’s scalability and the quality of its derived diversity metrics, emphasizing its potential to improve exploration and diversity in complex, open-ended optimization tasks. Source code is available on GitHub: https://github.com/ld-ing/qdhf.
@@ 1 Introduction
Foundation models such as large language models (LLMs) and text-to-image generation models in effect compress vast archives of human culture into powerful and flexible tools, serving as a foundation for diverse down-stream applications (Brown et al., 2020; Bommasani et al., 2021). Their promise includes helping individuals better meet their diverse goals, such as exploring their creativity in different modalities and coming up with novel solutions. One mechanism to build upon such foundational knowledge is reinforcement learning from human feedback (RLHF, Christiano et al. (2017)), which can make models both easier to use (by aligning them to human instructions), and more competent (by improving their capabilities based on human preferences).
@@
RLHF is a relatively new paradigm, and deployments of it often follow a relatively narrow recipe as maximizing a learned reward model of averaged human preferences over model responses. This work aims to broaden that recipe to include optimizing for interesting diversity among responses, which is of practical importance for many creative applications such as generative text-to-image models (Rombach et al., 2022), and may have the potential to further improve complex and open-ended learning tasks (through improved exploration), personalization (to serve individual rather than average human preference), and fairness (to embrace diversity and counter-balance algorithmic biases in gender, ethnicity and more).
@@
Diversity encourages exploration, which is essential for finding novel and effective solutions to complex problems. Without diversity, optimization algorithms might converge prematurely to suboptimal solutions, resulting in getting stuck in local optima or producing only a limited set of responses (i.e., mode collapse). The diversity aspect is especially substantial in quality diversity (QD) algorithms (Pugh et al., 2016; Cully et al., 2015; Lehman & Stanley, 2011b; Mouret & Clune, 2015), where diversity metrics are explicitly utilized to encourage and maintain the variation of solutions during optimization.
@@
The main idea is to derive distinct representations of what humans find interestingly different, and incorporate this procedure in diversity-driven optimization algorithms. We introduce a new concept, Quality Diversity through Human Feedback (QDHF), which empowers QD algorithms with diversity metrics actively learned from human feedback during optimization. To illustrate this concept, we propose an implementation of QDHF capable of formulating arbitrary diversity metrics using latent space projection, and aligning them with human feedback through contrastive learning (Hadsell et al., 2006; Dosovitskiy et al., 2014).
@@
Our work is inspired by the considerable benefits that the learned reward models have unlocked for RLHF. Analogous to reward functions, diversity metrics are often found qualitative, complex, and challenging to exhaustively specify. While existing QD algorithms demonstrate proficiency in addressing complex search and generative tasks due to their inherent explorative capabilities, their reliance on manually crafted diversity metrics restricts their applicability in real-world open-ended tasks. QDHF aims to bridge this gap, allowing QD algorithms to easily adapt to more challenging tasks by actively learning diversity metrics through iterative exploration and human feedback.
In summary, the main contributions of this work are: 1) Introducing QDHF and its implementation leveraging latent projection and contrastive learning. 2) Demonstrating that QDHF mirrors the search capabilities inherent in QD algorithms with manually crafted diversity metrics and outperforms comparable methods that employ automated diversity detection in benchmark robotics QD tasks. 3) Implementing QDHF in the latent space illumination (LSI) task for image generation, demonstrating its capability in producing diverse, high-quality images utilizing human feedback. 4) Providing an analysis on QDHF’s sample efficiency and the quality of its learned diversity metrics.
@@ 2 Preliminaries
In this section, we cover the basic and most relevant aspects of quality diversity (QD) algorithms (Mouret & Clune, 2015; Cully et al., 2015; Pugh et al., 2016; Lehman & Stanley, 2011b), and refer the readers Appendix A for more detailed descriptions.
@@ 2.1 Quality Diversity
QD algorithms effectively explore the search space by maintaining diverse high-quality solutions and using them to drive optimization. Mathematically, given a solution space $\mathcal{X}$, QD considers an objective function $J:\mathcal{X}\rightarrow\mathbb{R}$ and $k$ diversity metrics $M_{i}:\mathcal{X}\rightarrow\mathbb{R}$, $i=1,2,\cdots,k$. The diversity metrics jointly form a measurement space $M(\mathcal{X})$, which quantifies the diversity of samples. For each unique measurement in $M(\mathcal{X})$, the global objective $J^{*}$ of QD is to find a solution $x\in\mathcal{X}$ that has a maximum $J(x)$ among all solutions with the same diversity measurement.
Considering that the measurement space $M(\mathcal{X})$ is usually continuous, a QD algorithm will ultimately need to find and store an infinite number of solutions corresponding to each location in the measurement space. One common way to mitigate this is to discretize $M(\mathcal{X})$ into an archive of $s$ cells $\{C_{1},C_{2},\cdots,C_{s}\}$, which was introduced in MAP-Elites (Mouret & Clune, 2015; Cully et al., 2015) and has been widely adopted in its variants. The QD objective is thus approximated by relaxing the global objective to finding a set of solutions $\{x_{i}\},i\in\{1,...,s\}$, and each $x_{i}$ is the best solution for one unique cell $C_{i}$. This (approximated) $J^{*}$ can be thus formulated as:
@@
2.2 Quality Diversity without Predefined Diversity Metrics
Diversity maintenance in QD algorithms usually relies on manually designed diversity metrics to ensure a varied solution archive. However, such a requirement restricts the applicability of QD in more complex and open-ended domains, where the notion of diversity is likely to be abstract and qualitative. To address this, the concept of QD with automatic discovery of diversity has become popular in recent studies. Instead of using pre-defined diversity metrics, Cully (2019); Meyerson et al. (2016); Grillotti & Cully (2022) proposed AURORA, which utilizes unsupervised dimension reduction techniques to learn diversity metrics directly from the data.
However, one problem with these unsupervised methods is that the derived diversity metrics tend to capture the overall variance in the existing data, which may not align well with the diversity required for obtaining novel and superior solutions, making these methods uncapable in solving complex problems. As inspired by recent work in RLHF (Christiano et al., 2017) where reward models learned from human feedback are used to solve abstract and complicated optimization problems, we raise two central research questions: Can human feedback be employed to derive appropriate diversity metrics for QD that benefit optimization? And how does this approach compare to unsupervised or manually crafted metrics in terms of diversity maintenance and effectiveness of search?
To address these questions, we introduce a new paradigm of quality diversity through human feedback (QDHF), where the diversity metrics are learned from human judgment on similarity of the data. Using human feedback is not only more flexible than manually designing a diversity metric, but it also caters to more abstract and complex domains where defining a numeric metric of diversity is challenging. In general, QDHF endeavors to optimize for diverse solutions by leveraging diversity metrics that align with human intuition. In the following section, we introduce a specific way of implementing QDHF through latent space projection and contrastive learning.
@@ 3 Methods
In this section, we first introduce our formulation of arbitrary diversity metrics from a general modeling perspective using latent space projection, then describe how to align such metrics with human feedback through contrastive learning. Finally, we propose a quality diversity through human feedback (QDHF) framework that leverages human judgment for quality diversity algorithms as a semi-supervised optimization method.
3.1 Diversity Characterization through Latent Projection
Recent work (Cully, 2019; Meyerson et al., 2016; Grillotti & Cully, 2022) has shown that unsupervised dimensionality reduction algorithms such as principal component analysis (PCA) and auto-encoder (AE) can be utilized to automatically learn robot behavioral descriptors based on raw sensory data. In our framework, we first introduce a general concept of diversity characterization. Given data that contains information of diversity, we implement a latent projection, transforming it into a semantically meaningful latent space. More specifically, given an input vector $x$, we first (optionally) employ a feature extractor $f:\mathcal{X}\rightarrow\mathcal{Y}$ where $\mathcal{X}$ denotes the input space and $\mathcal{Y}$ the feature space. Post extraction, a projection function parameterized with $\theta$, denoted as $D_{r}(y,\theta):\mathcal{Y}\rightarrow\mathcal{Z}$, is applied to project the feature vector $y$ into a more compact latent representation:
@@ 3.1 Diversity Characterization through Latent Projection
In this context, $\mathcal{Z}$ represents the latent space, wherein each axis or dimension corresponds to a diversity metric derived from the data. The magnitude and direction along these axes capture nuanced characteristics of the original data, offering a compact yet informative representation of $x$. For example, on a single axis, smaller and larger values typically indicate variations in certain diversity characteristics. In this work, we apply linear projection for dimensionality reduction, where the parameters are learned through a contrastive learning process described in the following subsection.
3.2 Aligning Diversity Metrics with Human Preferences
While classic unsupervised dimensionality reduction methods capture significant data variances, the resulting latent representations may not consistently offer meaningful semantics for optimization. Recognizing this shortcoming, QDHF effectively aligns the diversity latent space with human notions of diversity by adopting contrastive learning Chopra et al. (2005); Schroff et al. (2015).
@@ 3.2 Aligning Diversity Metrics with Human Preferences
Recent work Radford et al. (2021); Chen et al. (2020); He et al. (2020); Tian et al. (2020) has explored using contrastive learning as a type of self-supervised representation learning strategy for images and natural language, and has demonstrated success of using it in modeling human preferences for RLHF (Christiano et al., 2017) and image perceptual similarity (Fu et al., 2023). Our framework takes a similar approach. Given a triplet of latent embeddings $z_{1}$, $z_{2}$, and $z_{3}$ and suppose that the human input indicates that $z_{1}$ is more similar to $z_{2}$ rather than $z_{3}$, our intention is to optimize the spatial relations of $z_{1}$ and $z_{2}$ in the latent space relative to $z_{3}$. We use a triplet loss mechanism, i.e., to minimize the distance between $z_{1}$ and $z_{2}$ while maximize between $z_{1}$ and $z_{3}$ via a hinge loss. This objective can be formalized as:
@@
where $D(\cdot,\cdot)$ represents a distance metric in the embedding space, and $m$ acts as a predetermined margin. Through this approach, the learned latent projections are aligned with both the inherent structure of the data and human notions of similarity.
Human judgement on similarity.
To accommodate our design of contrastive learning, we use the Two Alternative Forced Choice (2AFC) mechanism (Zhang et al., 2018; Fu et al., 2023) to obtain human judgments on input triplets. When presented with a triplet $\{x_{1},x_{2},x_{3}\}$, an evaluator is prompted to discern whether $x_{2}$ or $x_{3}$ is more similar to the reference, $x_{1}$. Importantly, this mechanism works not only with human judgment, but also with judgment produced by heuristics and other AI systems, meaning that our framework remains universally applicable across different feedback modalities.
@@ 3.3 Quality Diversity through Human Feedback
Based on the QDHF concept introduced in Sec. 2.2, we propose an implementation of QDHF through latent space projection and contrastive learning with human judgments.
@@
where $z=D_{r}(f(x),\theta)$ (Eq. 2). To effectively learn the parameters $\theta$ in latent projection, we utilize the contrastive learning mechanism (Eq. 3).
Training strategy.
One essential component in the QDHF framework is the effective collection and utilization of human feedback. A straightforward strategy is to initially sample randomly generated solutions, then collect human judgment on the similarity of these samples, and finally train the diversity representations. These representations can then be used as diversity metrics for QD algorithms in a standard way. We term this approach ‘QDHF-baseline’, which serves as a baseline method that utilizes human feedback in QD optimization.
@@
A potential drawback, however, of this baseline approach is that, for complex problems, the randomly generated solutions are often low-quality. As optimization proceeds, higher-quality solutions are found more frequently, but the diversity representations may not well-reflect the diversity between those. We propose a progressive training strategy for QDHF, which uses an active learning strategy such that the latent projection is initially trained with randomly sampled human judgment data, but also fine-tuned throughout the QD process. As QD continually adds new solutions to its archive, we select samples from this evolving archive to gather human judgments on their similarity. This online feedback is then used to iteratively fine-tune the latent projection. Subsequently, the QD algorithm is re-initialized with current solutions but integrated with updated diversity representations. This method ensures that the diversity representations remain relevant and reflective of the improved quality of solutions as optimization advances.
The frequency of fine-tuning decreases exponentially over time as the learned metrics become more robust. In our experiments, the latent projection is updated 4 times at iteration $1$, $10\%\cdot n$, $25\%\cdot n$, and $50\%\cdot n$ for a total of $n$ iterations. To fairly compare with QDHF-offline, each update consumes 1/4 of the total budget of human feedback.

### v3

@@ Abstract
Reinforcement Learning from Human Feedback (RLHF) has shown potential in qualitative tasks where easily defined performance measures are lacking. However, there are drawbacks when RLHF is commonly used to optimize for average human preferences, especially in generative tasks that demand diverse model responses. Meanwhile, Quality Diversity (QD) algorithms excel at identifying diverse and high-quality solutions but often rely on manually crafted diversity metrics. This paper introduces Quality Diversity through Human Feedback (QDHF), a novel approach that progressively infers diversity metrics from human judgments of similarity among solutions, thereby enhancing the applicability and effectiveness of QD algorithms in complex and open-ended domains. Empirical studies show that QDHF significantly outperforms state-of-the-art methods in automatic diversity discovery and matches the efficacy of QD with manually crafted diversity metrics on standard benchmarks in robotics and reinforcement learning. Notably, in open-ended generative tasks, QDHF substantially enhances the diversity of text-to-image generation from a diffusion model and is more favorably received in user studies. We conclude by analyzing QDHF’s scalability, robustness, and quality of derived diversity metrics, emphasizing its strength in open-ended optimization tasks. Code and tutorials are available at https://liding.info/qdhf.
Machine Learning, ICML
@@ 1 Introduction
Foundation models such as large language models (LLMs) and text-to-image generation models in effect compress vast archives of human knowledge into powerful and flexible tools, serving as a foundation for down-stream applications (Brown et al., 2020; Bommasani et al., 2021). Their promise includes helping individuals better meet their varied goals, such as exploring their creativity in different modalities and coming up with novel solutions. One mechanism to build upon such foundational knowledge is reinforcement learning from human feedback (RLHF) (Christiano et al., 2017), which can make models both easier to use (by aligning them to human instructions), and more competent (by improving their capabilities based on human preferences).
@@
RLHF is a relatively new paradigm, and its deployments often follow the relatively narrow recipe of maximizing a learned reward model of averaged human preferences over model responses. This work aims to broaden that recipe to include optimizing for interesting diversity among responses, which is of practical importance for many creative applications such as text-to-image generation (Rombach et al., 2022). Such increased diversity can also improve optimization for complex and open-ended tasks (through improved exploration), personalization (serving individual rather than average human preference), and fairness (to offset algorithmic biases in gender, ethnicity, and more).
@@
Diversity encourages exploration, which is essential for finding novel and effective solutions to many complex problems. Without diversity, optimization algorithms can converge prematurely, resulting in getting stuck in local optima or producing only a limited set of responses (i.e., mode collapse). Diversity-seeking is a core aspect of Quality Diversity (QD) algorithms (Pugh et al., 2016; Cully et al., 2015; Lehman & Stanley, 2011b; Mouret & Clune, 2015), where diversity metrics are explicitly defined and utilized to encourage the variation of solutions during optimization.
@@
Our main idea is to derive distinct representations of what humans find interestingly different and use such diversity representations to support optimization. We introduce a novel method, Quality Diversity through Human Feedback (QDHF), which empowers QD algorithms with diversity metrics actively learned from human feedback during optimization. QDHF serves as an online, diversity-driven optimizer for open-ended tasks. It navigates the search space through the application of progressively learned diversity, which is achieved by concurrently fine-tuning inferred diversity by aligning it to human feedback in response to the discovery of novel solutions.
To illustrate its capability, we propose a generic implementation of QDHF, which is capable of formulating arbitrary diversity metrics using latent space projection, and aligning them to human feedback through contrastive learning (Hadsell et al., 2006; Dosovitskiy et al., 2014). Experiments are conducted on benchmarks across three domains: robotics, reinforcement learning (RL), and generative modeling, demonstrating QDHF’s superior performance in producing diverse, high-quality responses.
@@
This work is inspired by the considerable benefits that learned reward models have unlocked for RLHF. Analogous to reward functions, diversity metrics are often qualitative, complex, and hard to specify. Existing QD algorithms demonstrate proficiency in addressing complex search tasks, but their reliance on manually crafted diversity metrics restricts their applicability in real-world open-ended tasks. QDHF aims to bridge this gap, allowing QD algorithms to easily adapt to more challenging tasks by learning diversity metrics through iterative exploration and alignment. In summary, the main contributions of this work are:
  1. 1.
    Introducing QDHF and its implementation leveraging latent projection and contrastive learning.
  2. 2.
    Demonstrating that QDHF mirrors the search capabilities of QD algorithms with manually crafted diversity metrics in benchmark robotics and RL tasks.
  3. 3.
    Implementing QDHF in the latent space illumination (LSI) task for open-ended text-to-image generation, showing its capability in improving diffusion model to generate more diverse and preferable responses.
  4. 4.
    Ablation studies on QDHF’s sample efficiency, robustness, and the quality of its learned diversity metrics.
@@ 2 Preliminaries
This section covers the basic and most relevant aspects of QD algorithms. More detailed descriptions are included in Appendix A.
@@ 2.1 Quality Diversity
QD algorithms effectively explore the search space by maintaining diverse high-quality solutions and using them to drive optimization. Given a solution space $\mathcal{X}$, QD considers an objective function $J:\mathcal{X}\rightarrow\mathbb{R}$ and $k$ diversity metrics $M_{i}:\mathcal{X}\rightarrow\mathbb{R}$, $i=1,2,\cdots,k$. The diversity metrics jointly form a measurement space $M(\mathcal{X})$, which quantifies the diversity of samples. For each unique measurement in $M(\mathcal{X})$, the global objective $J^{*}$ of QD is to find a solution $x\in\mathcal{X}$ that has a maximum $J(x)$ among all solutions with the same diversity measurement.
Considering that the measurement space $M(\mathcal{X})$ is usually continuous, a QD algorithm will ultimately need to store an infinite number of solutions corresponding to each solution in the measurement space. One common way to mitigate this is to discretize $M(\mathcal{X})$ into an archive of $s$ cells $\{C_{1},C_{2},\cdots,C_{s}\}$, which was introduced in MAP-Elites (Mouret & Clune, 2015; Cully et al., 2015) and has been widely adopted. The QD objective is thus approximated by a relaxation to finding a set of solutions $\{x_{i}\},i\in\{1,...,s\}$, and each $x_{i}$ is the best solution for one unique cell $C_{i}$. This (approximated) $J^{*}$ is:
@@
2.2 Quality Diversity w/o Predefined Diversity Metrics
Diversity maintenance in QD usually relies on manually designed diversity metrics to ensure a varied solution archive. However, such a requirement restricts the applicability of QD in more complex and open-ended domains, where the notion of diversity is likely to be abstract and qualitative. To address this, QD with automatic discovery of diversity has become popular. Instead of using pre-defined diversity metrics, recent work such as AURORA (Cully, 2019; Grillotti & Cully, 2022b) and RUDA (Grillotti & Cully, 2022a) utilize unsupervised dimension reduction techniques to learn diversity metrics directly from the data.
One limitation of unsupervised methods is that the derived diversity metrics often capture the overall variance in current solutions, which may not align well with the diversity needed for discovering novel and superior solutions. Such misalignment requires explicit mechanisms (e.g., RUDA) to guide the search, making the methods less applicable in complex and open-ended environments. We introduce the new paradigm of QDHF, which offers greater flexibility compared to manually designing diversity metrics and outperforms unsupervised methods by leveraging diversity metrics that align to human intuition.
@@ 3 Methods
This section first introduces a formulation of arbitrary diversity metrics from a generic learning perspective using latent projection, then describes how to align such metrics to human intuition of diversity using contrastive learning. Finally, we introduce quality diversity through human feedback (QDHF), an online method for open-ended diversity-driven optimization. An overview is shown in Fig. 1.
Figure 1: Overview of the proposed QDHF method.
3.1 Diversity Characterization with Latent Projection
Prior work (Meyerson et al., 2016; Cully, 2019; Grillotti & Cully, 2022b) has shown that unsupervised dimensionality reduction algorithms such as principal component analysis (PCA) and auto-encoders (AE) can be utilized to automatically learn robot behavioral descriptors based on sensory data. In our framework, we first introduce a concept of diversity characterization, which uses latent projection to transform arbitrary data into a semantically meaningful latent space that represents diversity. More specifically, given an input vector $x$, we first employ a feature extractor $f:\mathcal{X}\rightarrow\mathcal{Y}$ where $\mathcal{X}$ denotes the input space and $\mathcal{Y}$ the feature space. Post extraction, a projection function parameterized with $\theta$, denoted as $D_{r}(y,\theta):\mathcal{Y}\rightarrow\mathcal{Z}$, is applied to project $y$ into a more compact latent representation:
@@ 3.1 Diversity Characterization with Latent Projection
$\mathcal{Z}$ represents the latent space, wherein each dimension corresponds to a diversity metric. The magnitude and direction capture various notions of diversity, offering a compact yet informative representation of $x$. For example, on a single dimension, smaller and larger values could indicate variations in certain characteristics such as the size of the object. In this work, we use linear projection for simplicity and generality, although non-linear multi-layer projections are also compatible with our framework. The parameters for the projection are learned through a contrastive learning process described in the following section.
3.2 Aligning Diversity Metrics to Human Intuition
While unsupervised dimensionality reduction methods capture significant data variances, the resulting latents may not consistently offer useful semantics for optimization. Recognizing this shortcoming, QDHF effectively aligns the latents to human notions of diversity through contrastive learning (Chopra et al., 2005; Schroff et al., 2015).
@@ 3.2 Aligning Diversity Metrics to Human Intuition
Recent work has explored using contrastive learning as a type of self-supervised representation learning strategy (Radford et al., 2021; Chen et al., 2020; He et al., 2020; Tian et al., 2020), and has demonstrated success of using it in modeling human preferences in RLHF (Christiano et al., 2017) and image similarity (Fu et al., 2023). Our framework takes a similar approach. Given a triplet of latent embeddings $z_{1}$, $z_{2}$, and $z_{3}$, and supposing that the human input indicates that $z_{1}$ is more similar to $z_{2}$ rather than $z_{3}$, our intention is to optimize the spatial relations of $z_{1}$ and $z_{2}$ in the latent space relative to $z_{3}$. We use a triplet loss mechanism, i.e., to minimize the distance between $z_{1}$ and $z_{2}$ while maximizing between $z_{1}$ and $z_{3}$ via a hinge loss. This objective can be formalized as:
@@
where $D(\cdot,\cdot)$ represents a distance metric in the embedding space, and $m$ acts as a predetermined margin. Through this approach, the latent projections are aligned to both the inherent structure of data and human notions of similarity.
Human judgment on similarity.
To accommodate our design, we use Two Alternative Forced Choice (2AFC) (Zhang et al., 2018; Fu et al., 2023) to obtain human judgments on the similarity of solutions. When presented with a triplet $\{x_{1},x_{2},x_{3}\}$, an evaluator is prompted to discern whether $x_{2}$ or $x_{3}$ is more similar to the reference solution, $x_{1}$. Importantly, this mechanism works not only with human judgment but also with judgment produced by heuristics and AI systems, meaning that our framework remains universally applicable across different feedback modalities.
@@ 3.3 Quality Diversity through Human Feedback
We hereby introduce QDHF and its implementation through latent space projection and contrastive learning.
@@
where $z=D_{r}(f(x),\theta)$ (Eq. 2). To effectively learn $\theta$, we use the contrastive learning mechanism (Eq. 3).
Training.
One essential component of QDHF is the effective utilization of human feedback. The initial diversity representations are learned from feedback collected on randomly generated solutions, which are often low-quality. But as optimization proceeds, novel and higher-quality solutions are found, and the initial diversity metrics are unlikely to characterize them correctly.
@@
We propose a progressive training strategy for QDHF where the latent projection is iteratively fine-tuned throughout the QD process. As QD continually adds new solutions to its archive, we select samples from this evolving archive to gather human judgments on their similarity, producing online feedback used to fine-tune the latent projection. Subsequently, the QD algorithm is re-initialized with current solutions but integrated with updated diversity representations. This method ensures that the diversity representations remain relevant and reflective of the improved quality of solutions as optimization advances, supporting continuous and open-ended optimization for complex tasks.
Baseline method.
We also propose a strong baseline method, ‘QDHF-Base,’ which integrates the general design principles of RLHF and QD algorithms but learns diversity metrics offline. This method first collects human judgments on the similarity of randomly generated solutions, then learns the diversity representations (akin to reward model learning in RLHF), and finally uses these representations as diversity metrics in QD. QDHF-Base, although a baseline, is a novel algorithm for combining RLHF with QD optimization, leveraging the proven effectiveness of the RLHF framework.
4 Theoretical Analysis
In this section, we include further interpretations of QDHF, provide theoretical backing, and relate the advantages of QDHF to prior work in online learning.
4.1 Learning Distinct Diversity Representations through Information Bottleneck
When the dimensionality of the feature space, particularly in the last layer of a network, is significantly reduced, the process compels the network to encapsulate essential discriminative information within a limited set of features. This process is effectively framed by the Information Bottleneck (IB) principle (Tishby et al., 2000), which suggests that an optimal representation, $Z$, of input data $X$, should conserve only the information necessary for predicting the output $Y$. The IB principle articulates this as an optimization task, aiming to balance the mutual information between $Z$ and $Y$ while constraining the information between $X$ and $Z$. Mathematically, this balance is captured by:
$\mathcal{L}_{IB}=I(X;Z)-\beta I(Z;Y).$ (5)
Here, a Lagrange multiplier $\beta$ serves as a trade-off parameter, moderating the amount of information about $X$ retained in $Z$ and ensuring the relevance of $Z$ for predicting $Y$. More recent work (Alemi et al., 2017) has further shown the IB principle’s pivotal role in deep neural networks’ capacity to compress and selectively filter information.
In the context of contrastive learning, the objective is to optimize representations that carry the necessary information for predicting human preferences for similarity. Conventional approaches often employ large embedding dimensions to maximize the capability for similarity estimation, such as embeddings with 512 dimensions in DreamSim (Fu et al., 2023). In contrast, QDHF opts for significantly reduced dimensions to produce compact representations, imposing a constraint on the mutual information between $Z$ and $X$, formulated by:
$I(X;Z)\leq I_{c}$ (6)
where $I_{c}$ is the information constraint. Intuitively, $I_{c}$ forces $Z$ to learn a minimal sufficient statistic of $X$ for predicting $Y$, resulting in $Z$ being concise yet information-rich representations. The IB principle suggests that such representations should be both discriminative and succinct if $\mathcal{L}_{IB}$ is minimal. Our empirical analysis, as described in Sec. 5.4, further validates that when $I(Z;Y)$ the accuracy of predicting human judgments on the validation set - is high, QDHF can learn robust and informative diversity metrics and closely approaches the capability of QD using ground truth diversity.
Complementing this, the contrastive loss function (Eq. 3) minimizes the distance between similar pairs and maximizes it between dissimilar pairs, thereby encouraging the learning of meaningful representations for diversity that are aligned with human intuition. These representations can thus serve as robust diversity metrics in QD optimization, enhancing the model’s generalization and creativity.
4.2 Quality Diversity as an Active Sampling Strategy for Online Learning
We also provide an analysis of QDHF from the perspective of online learning. Prior work has explored using online learning for similarity, especially in large-scale data environments. QDHF introduces a nuanced methodology, enhancing traditional models like the Online Algorithm for Scalable Image Similarity (OASIS) (Chechik et al., 2010). OASIS, part of the Passive-Aggressive (PA) (Crammer et al., 2006) algorithm family, optimizes similarity across image triplets $(p_{i},p_{i}^{+},p_{i}^{-})$ to ensure a predefined margin in parameterized similarity scores $S_{W}$, formalized as:
$l_{W}(p_{i},p_{i}^{+},p_{i}^{-})=\max\left(0,1-S_{W}(p_{i},p_{i}^{+})+S_{W}(p_{i},p_{i}^{-})\right).$ (7)
This contrastive loss (equivalent to Eq. 3 in the main paper by having a margin of $1$) leads to the global loss $L_{W}$ defined as the sum of hinge losses over all triplets. The loss is minimized through iterative PA updates, where each iteration $i$ involves optimizing $W$ via:
[…]
subject to $l_{W}(p_{i},p_{i}^{+},p_{i}^{-})\leq\xi$ and $\xi\geq 0$, where $\|\cdot\|_{F}$ denotes the Frobenius norm. The updates make sure that $W_{i}$ optimizes the balance between staying close to the previous parameters $W_{i-1}$ and reducing loss on the current triplet $l_{W}(p_{i},p_{i}^{+},p_{i}^{-})$, with $C$ modulating this trade-off.
Conversely, QDHF focuses on QD as its core mechanism, utilizing contrastive learning to derive representations that serve as both benchmarks for diversity and guides for exploration. Unlike OASIS’s fixed-margin strategy, QDHF employs an adaptive scheme that refines its diversity metrics through continuous feedback and new data integration, fostering a progressing representation space that captures complex diversity-exploration interrelations. Mathematically, QDHF seeks to minimize the global loss as a blend of sustained learning from historical preference data and online data with feedback based on new triplet samplings:
$L_{W}=\sum_{\text{old triplets}}l_{W}+\sum_{\text{new triplets}}l_{W},$ (9)
where $\sum_{\text{old triplets}}l_{W}$ aims to retain the learned diversity, similar to how the first term $\|W-W_{i-1}\|_{F}^{2}$ in Eq. 8 ensures continuity. Theoretical analysis and empirical evidence from past studies (Crammer et al., 2006) suggest that such iterative online algorithms can achieve a cumulative online loss that remains small and converges over time.
While OASIS leverages a structured update rule to ensure that the learned similarity matrix progressively aligns with the target function, QDHF employs continuous feedback loops to develop a learning environment for diversity and exploration. This strategic feedback mechanism can be viewed as using QD as a sampling strategy for actively discovering unseen data for refining the learned diversity metrics, which ensures a vibrant, adaptive optimization landscape for online learning in more complex, open-ended environments.